# Orthogonal Blocking Kaczmarz Algorithm Based on Preprocessing Technology

### v2

@@ Abstract
In this paper, a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques is proposed to solve large-scale sparse linear systems $Ax=f$. Firstly, the Reverse Cuthill McKee Algorithm (RCM) algorithm is used to preprocess the linear system, and then a new partitioning strategy is proposed to divide orthogonal blocks into one category, which greatly reduces computational complexity. In addition, the convergence of POBK algorithm is theoretically proven, and a corresponding analysis of its faster convergence is also provided. Finally, numerical simulations indicate that our method outperforms the latest Kaczmarz methods such as GRBK, RBK(k) and GREBK(k) in both iteration steps (IT) and CPU time aspects.
@@
Kaczmarz algorithm, Preprocessing techniques, Orthogonal partitioning, Robustness, k-means
@@ 1 Introduction
This paper mainly considers the solution of a large uniformly sparse linear system in the form of
@@
where $A\in{{\mathbb{R}}^{m\times m}}$, $f\in{{\mathbb{R}}^{m}}$, and $x\in{\mathbb{R}^{m}}$. One of the classic and effective iterative projection methods for solving such problems is the Kaczmarz method.
The classic Kaczmarz method scans each row of the matrix in a cyclic manner and projects its current iterative solution $x_{j}$ onto a hyperplane $H_{i}=\{x\mid\left\langle A_{i},x\right\rangle=f_{i}\}$ defined by the system rows 1 ; 2 ; 19 . Specifically, the initial value is set to $x_{0}$, the iteration format of the classic Kaczmarz method is
$x_{j+1}=x_{j}+\frac{f_{i}-A_{i}^{T}x}{\left\|A_{i}\right\|_{2}^{2}}\left(A_{i}^{T}\right).$ (1.2)
The convergence of the classic Kaczmarz method depends on the order of matrix row selection, and the convergence speed is slow and it is difficult to provide a convergence theory. Therefore, Strohmer and Vershynin 3 ; 4 proposed a random Kaczmarz (RK) algorithm with an expected exponential convergence rate. However, the probability criterion for randomly selecting rows in the RK algorithm is not optimal. Based on the idea of greed, Ansorge proposed the greedy residual Kaczmarz (GEK) method 5 . Nutini et al. proposed the greedy distance Kaczmarz (GDK) method 6 . Bai Zhongzhi and Wu Wenting combined the mathematical concept of randomness and introduced a new and more effective selection criterion by defining an indicator set. They proposed the greedy random Kaczmarz (GRK) algorithm and its relaxed version, which can converge to the system’s solution faster. Liu et al. proposed a variant of GRK to solve the ridge regression problem 9 . Du Yishu et al. introduced an effective distance based row selection criterion and proposed the distance based greedy random Kaczmarz (GDRK) algorithm 10 , which has better solving performance than traditional RK and GRK algorithms.
The block Kaczmarz (BK) method 11 ; 12 is a natural extension of the classical Kaczmarz method and another effective way to improve its convergence performance. Unlike the row Kaczmarz method, the block Kaczmarz algorithm involves selecting multiple row metrics in each iteration. Given an initial solution $x_{0}$, in the k-th iteration, first select a subset of travel indicators $\tau_{j}\in\left[m\right]$, and then project the iterative solution $x_{j-1}$ onto the solution space $A_{\tau_{j}}x=f_{\tau_{j}}$. Using the pseudo inverse of $A_{\tau_{j}}$, the Kaczmarz iteration formula for the block is
@@
where $A_{\tau_{j}}$ and $f_{\tau_{j}}$ respectively represent the submatrix and right-hand vector corresponding to $A$ and $f$. Needell et al. proposed the Good Saving and Block Random Kaczmarz (RBK) algorithms 13 ; 14 ; 15 . Subsequent scholars combined RBK with greedy algorithms and proposed the Kaczmarz algorithm for greedy random partitioning (GRBK) 18 . Zhang et al. proposed the greedy random Motzkin Caczmarz (GRMK) method using the greedy selection rule on the maximum residual 16 . Based on the k-means clustering algorithm, Jiang, Li et al. proposed the row clustering block Kaczmarz (RBK (k)) algorithm based on the k-means algorithm 17 and the global random block Kaczmarz method 18 , further improving the convergence rate of the algorithm. On this basis, Zheng Wenhao et al. explored the relationship between ordinary residuals and standardized residuals, and constructed a new random greedy residual block Kaczmarz (GREBK (k)) algorithm 20 , which has a much faster computational speed than the RBK (k) method.
@@
In this article, inspired by the GRBK, RBK(k), and GREBK(k) algorithms, the convergence rate of the block iteration Kaczmarz algorithm is indeed much faster than that of the row iteration Kaczmarz algorithm. However, our analysis shows that the sparse structure of the matrix still has a significant impact on the convergence rate of the block Kaczmarz algorithm. When faced with sparsely distributed matrices, the iteration time for the block Kaczmarz algorithm to obtain accurate solutions is not ideal. To improve the convergence speed of block Kaczmarz and the convergence of iterative sequences, this paper constructs an orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques. The algorithm mainly uses preprocessing techniques to rearrange the original matrix, making the new matrix an approximate block diagonal matrix. Then, the matrix is divided into blocks, with the aim of grouping orthogonal blocks into one category. In this way, each iteration in each category undergoes two orthogonal projections on two mutually orthogonal hyperplanes, so the iterative solution can approach the exact solution more quickly. This article theoretically proves the convergence of the POBK algorithm. The numerical experimental results also show that compared with the GRBK, RBK(k), and GREBK(k) algorithms, the POBK algorithm is more efficient and ideal in terms of iterative sequence convergence and algorithm running time.
@@
The structure of this article is as follows. In Section 2, we briefly introduced the GRBK algorithm, RBK (k) algorithm, and GREBK (k) algorithm. In Section 3, we introduced the preprocessing techniques and partitioning ideas used, analyzed the reasons for the selection, and further proposed a variant POBK algorithm of the block Kaczmarz algorithm. In Section 4, we proved the convergence of the proposed method. In Section 5, we selected appropriate matrices for numerical simulation to demonstrate the effectiveness of our method. Finally, in Section 6, we provide conclusions and point out some future directions for work.
@@ 2 Introduction to existing block methods
The emergence of the block Kaczmarz method has greatly improved the running time and convergence rate of the Kaczmarz method in handling high-dimensional linear systems. The block Kaczmarz method under greedy algorithms and random strategies is a natural development of the block Kaczmarz method. The specific description of the algorithm can be found in Algorithm 1.
Algorithm 1 GRBK Algorithm 17
@@
3:Randomly block $A$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ for row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
@@
In 2022, based on the principle of block Kaczmarz method, Jiang Xianglong et al. first combined k-means clustering algorithm with greedy random technique to propose a novel random block Kaczarz (RBK (k)) method 17 . Based on the distance between the hyperplane formed by each row of the sparse matrix and the current iteration solution, k-means clustering is performed to divide the rows corresponding to the hyperplane with similar distances into one class. In each iteration, the class with the largest distance is selected for orthogonal projection, and the practical probability criterion for selecting the working block sub-matrix in each iteration is considered an effective variant of the relaxed greedy random Kaczmarz algorithm. The specific description of the RBK(k) algorithm can be found in Algorithm 2.
Algorithm 2 RBK(k) Algorithm 17
@@
In 2023, Zheng Wenhao, et al. discovered the relationship between ordinary residuals and standardized residuals 20 , that is, the standardized residuals are the 2-norm of the row vector of the ordinary residuals divided by the corresponding sparse matrix $A$
$\hat{r}_{i}=\frac{r_{i}}{\left\|A_{i}^{T}\right\|_{2}}=\frac{f_{i}-A_{i}^{T}x_{0}}{\left\|A_{i}^{T}\right\|_{2}}.$ (2.1)
@@
and the absolute value of standardized residuals is also a measure of greed in the mainstream greedy Kaczmarz algorithm. So they changed the k-means clustering objects in the RBK(k) algorithm from matrix $A$ to standardized residuals, and obtained the blocks of $A$. They proposed a block Kaczmarz (GREBK(k)) algorithm that combines the k-means algorithm to block the standardized residuals clustering, as shown in Algorithm 3.
Algorithm 3 GREBK(k) Algorithm 20
@@
3 Orthogonal Blocking Kaczmarz Method Based on Preprocessing Technology
3.1 Preprocessing Technology–Reverse Cuthill McKee Algorithm
As is well known, the Kaczmarz algorithm has good convergence for linear systems, especially for sparse linear systems. However, Kaczmarz algorithm does not exhibit good convergence speed for all sparse matrices, especially for matrices with large dimensions or scattered structural distributions. Even using block clustering and greedy algorithms cannot effectively change this situation. We analyze the iterative characteristics of the Kaczmarz algorithm and the structure of matrices in linear systems and find that due to Kaczmarz’s orthogonal projection onto the hyperplane during each iteration, the convergence speed and the angle between the hyperplane where the iteration point is located and the next projection hyperplane are closely related. Obviously, the smaller the angle between the two hyperplanes, the more iterations, and the slower the convergence. The closer the angle is to the perpendicular, the fewer iterations, and the faster the convergence. Taking a two-dimensional linear system as an example, Figure 1 shows the iterative differences between the two cases. $H_{1}$, $H_{2}$, $H_{3}$ represent the three hyperplanes of the linear system, $x_{0}$ represents the initial point, $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ represent the results of Kaczmarz’s first, second, third, and fourth iterations, respectively. The red dots where the hyperplanes intersect represent the exact solution $x_{*}$. The left side of Figure 1 represents the iteration process with a smaller angle between $H_{1}$ and $H_{2}$, and the right side represents the iteration process when $H_{2}$ is perpendicular to $H_{3}$, It can be clearly seen that when two hyperplanes are orthogonal to each other, the Kaczmarz algorithm only needs two steps to reach the exact solution, while the non orthogonal hyperplanes at this time are still far from the exact solution even after four steps of Kaczmarz iteration. So, we consider whether there is a method to extract rows that are orthogonal to each other in a linear system, or to process the original matrix into a matrix with rows that are orthogonal to each other? If so, the iteration time of the Kaczmarz algorithm will be greatly reduced, see 21 .
@@ 3.1 Preprocessing Technology–Reverse Cuthill McKee Algorithm
The Reverse Cuthill McKee (RCM) algorithm was initially used for reordering symmetric matrices. It will generate a matrix with rows and columns numbered according to the node numbers. By appropriately renumbering nodes, a new matrix with much smaller matrix bandwidth can usually be generated. The specific principle is that for a symmetric sparse matrix $A$ with a given bandwidth, the RCM algorithm can provide a permutation matrix $P$ such that $\widetilde{A}=PAP^{T}$ has a smaller bandwidth, i.e., $\widetilde{A}$ has diagonal elements closer to the diagonal.
Fortunately, the RCM algorithm also performs equally well on asymmetric matrices. So, in our proposed algorithm, we first preprocess the linear system using the RCM algorithm, transforming the original coefficient starting matrix into an approximate (block) diagonal matrix, see Figure 2. Obviously, there are many rows or row blocks that are mutually orthogonal in the approximate (block) diagonal matrix, see Figure 3. And this improves the drawback of the original matrix structure being arranged randomly. According to our above analysis, using the Kaczmarz algorithm again will greatly accelerate the convergence speed.
@@
3.2 Proposed methods

### v3

# Orthogonal Block Kaczmarz Algorithm Based on Preprocessing Technology

@@ Abstract
In this paper, a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques is proposed to solve large-scale sparse linear systems $Ax=f$. Firstly, the sparse Reverse Cuthill-McKee Algorithm (RCM) is used to preprocess the linear system, and then a new partitioning strategy is proposed to divide the mutually orthogonal blocks into one category, which greatly reduces computational complexity. In addition, we provide a theoretical proof of the convergence of the POBK algorithm and analyze its faster convergence. Finally, numerical simulations demonstrate that our method outperforms the latest Kaczmarz methods such as GRBK, RBK(k), GREBK(k), and aRBK in terms of CPU time.
@@
Kaczmarz algorithm, Sparse Reverse Cuthill-McKee ordering, Preprocessing techniques, Orthogonal partitioning
@@ 1 Introduction
This paper mainly considers the solution of a large sparse linear system in the form of
@@
where $A\in{{\mathbb{R}}^{m\times m}}$, $f\in{{\mathbb{R}}^{m}}$, and $x\in{\mathbb{R}^{m}}$. Nowadays, one of the classic and effective iterative projection methods for solving such problems is the Kaczmarz method [1, 2].
The classic Kaczmarz method scans each row of the matrix in a cyclic manner and projects its current iterative solution $x_{j}$ onto a hyperplane $H_{i_{k}}=\{x\in{\mathbb{R}^{m}}\mid A_{(i_{k})}x=f_{i_{k}}\}$ defined by the system rows [1, 2, 25]. Specifically, the initial value is set to $x_{0}$, assuming that the $i_{k}$th row has been selected at the $j$th iteration, then the $(j+1)$th estimate vector $x_{j+1}$ is obtained by:
$x_{j+1}=x_{j}+\frac{f_{i_{k}}-A_{(i_{k})}x_{j}}{\left\|A_{(i_{k})}\right\|_{2}^{2}}A_{(i_{k})}^{T}.$ (1.2)
The convergence of the classic Kaczmarz method depends on the order of matrix row selection, and the convergence speed is slow and it is difficult to provide a convergence theory. Therefore, Strohmer and Vershynin [3, 4] proposed a random Kaczmarz (RK) algorithm with an expected exponential convergence rate. However, the probability criterion for randomly selecting rows in the RK algorithm is not optimal. Based on the idea of greed, Ansorge proposed the greedy residual Kaczmarz (GEK) method [5]. Nutini et al. proposed the greedy distance Kaczmarz (GDK) method [6]. Bai Zhongzhi and Wu Wenting combined the mathematical concept of randomness and introduced a new and more effective selection criterion by defining an indicator set. They proposed the greedy random Kaczmarz (GRK) algorithm and its relaxed version [7, 8], which can converge to the system’s solution faster. Liu et al. proposed a variant of GRK to solve the ridge regression problem [9]. Du Yishu et al. introduced an effective distance-based row selection criterion and proposed the distance-based greedy random Kaczmarz (GDRK) algorithm [10], which has better solving performance than traditional RK and GRK algorithms.
The block Kaczmarz (BK) method [11, 12] is a natural extension of the classical Kaczmarz method and provides another effective way to improve its convergence performance. Unlike the row Kaczmarz method, the block Kaczmarz algorithm involves selecting multiple row indices in each iteration. Given an initial solution $x_{0}$, in the $k$-th iteration, first select a subset of row indices $\tau_{j}\in\left[m\right]$, and then project the iterative solution $x_{j-1}$ onto the solution space $A_{\tau_{j}}x=f_{\tau_{j}}$. Using the pseudo-inverse of $A_{\tau_{j}}$, the Kaczmarz iteration formula for the block is
@@
where $A_{\tau_{j}}$ and $f_{\tau_{j}}$ respectively represent the submatrix and right-hand vector corresponding to $A$ and $f$.
@@
Recently, Needell et al. proposed the good saving and block random Kaczmarz (RBK) algorithms [13, 14, 15]. Necoara developed a family of randomized block Kaczmarz algorithms that use a subset of constrained and extrapolated steps at each step, and can be deployed on distributed computing units [16]. Subsequent scholars combined RBK with greedy algorithms and proposed the Kaczmarz algorithm for greedy random partitioning (GRBK) [19]. Zhang et al. proposed the greedy random Motzkin Caczmarz (GRMK) method using the greedy selection rule on the maximum residual [17]. Based on the k-means clustering algorithm, Jiang, Li et al. proposed the row clustering block Kaczmarz (RBK(k)) algorithm based on the k-means algorithm [18] and the global random block Kaczmarz method [19], further improving the convergence rate of the algorithm. On this basis, Zheng et al. explored the relationship between ordinary residuals and standardized residuals, and constructed a new random greedy residual block Kaczmarz (GREBK(k)) algorithm [20], which has a much faster computational speed than the RBK(k) method. In 2024, Zhang et al. combined fast projection and weight averaging techniques to propose the aRBK algorithm [21].
@@
In this article, inspired by the GRBK, RBK(k), and GREBK(k) algorithms, the convergence rate of the block iteration Kaczmarz algorithm is indeed much faster than that of the row iteration Kaczmarz algorithm. However, our analysis shows that the sparse structure of the matrix still has a significant impact on the convergence rate of the block Kaczmarz algorithm. For the sparse matrix with a scattered distribution, the iteration time for obtaining an exact solution using the block Kaczmarz algorithm is not ideal. To improve the convergence speed of block Kaczmarz, this paper constructs a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing technology. The proposed algorithm utilizes preprocessing techniques to rearrange the original matrix so that it becomes approximately a banded matrix. Then, the matrix is divided into several blocks, with the goal of grouping orthogonal blocks into one category. In this way, each iteration in this category is performed as two orthogonal projections on nearly orthogonal hyperplanes, so that the iterative solution converges faster to the exact solution. This article theoretically proves the convergence of the POBK algorithm. The numerical experimental results also show that compared with the GRBK, RBK(k), GREBK(k) and aRBK algorithms, the POBK algorithm is more efficient and ideal in terms of algorithm running time.
@@
The structure of this article is as follows: In Section 2, we briefly introduced the GRBK, RBK(k), GREBK(k) and aRBK algorithms. In Section 3, we introduce the preprocessing techniques and partitioning ideas used, analyzed the reasons for their selection, and further proposed a variant POBK algorithm of the block Kaczmarz algorithm. In Section 4, we proved the convergence of the proposed method. In Section 5, we selected some matrices for numerical simulation to demonstrate the effectiveness of our method. Finally, in Section 6, we provide conclusions and point out some future directions for work.
@@ 2 Introduction to existing block methods
The emergence of the block Kaczmarz method has greatly improved the running time and convergence rate of the Kaczmarz method in handling high-dimensional linear systems. The block Kaczmarz method under greedy algorithms and random strategies (GRBK) is a natural development of the block Kaczmarz method. The specific description of the algorithm can be found in Algorithm 1.
Algorithm 1 GRBK Algorithm [18]
@@
3:Randomly obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
@@
In 2022, based on the principle of block Kaczmarz method, Jiang Xianglong et al. first combined k-means clustering algorithm with greedy random technique to propose a novel random block Kaczarz (RBK(k)) method [18]. Based on the distance between the hyperplane formed by each row of the sparse matrix and the current iteration solution, k-means clustering is performed to divide the rows corresponding to the hyperplane with similar distances into one class. In each iteration, the class with the largest distance is selected for orthogonal projection, and the practical probability criterion for selecting the working block submatrix in each iteration is considered an effective variant of the relaxed greedy random Kaczmarz algorithm. The specific description of the RBK(k) algorithm can be found in Algorithm 2.
Algorithm 2 RBK(k) Algorithm [18]
@@
In 2024, Zheng Wenhao, et al. discovered the relationship between ordinary residuals and standardized residuals [20], that is, the standardized residuals are the 2-norm of the row vector of the corresponding sparse matrix $A$ divided by the ordinary residuals.
$\hat{r}_{i}=\frac{r_{i}}{\left\|A_{(i)}\right\|_{2}}=\frac{f_{i}-A_{(i)}x_{0}}{\left\|A_{(i)}\right\|_{2}}.$ (2.1)
@@
and the absolute value of standardized residuals is also a measure of greed in the mainstream greedy Kaczmarz algorithm. So they changed the k-means clustering objects in the RBK(k) algorithm from matrix $A$ to standardized residuals, and obtained the partitioning strategy for matrix $A$. They proposed a block Kaczmarz (GREBK(k)) algorithm that combines the k-means algorithm to the block standardized residuals clustering, as shown in Algorithm 3.
Algorithm 3 GREBK(k) Algorithm [20]
@@
In 2024, Zhang et al. [21] proposed a variant of the random Kaczmarz algorithm, the aRBK algorithm, by combining block projection and weight averaging techniques. Their combination can balance convergence speed, convergence range, and computational complexity. Please refer to Algorithm 4 for more details.
Algorithm 4 aRBK Algorithm [21]
1:$A$, $f$, $x_{0}$, $\left\{\omega_{j}\right\}_{j\in\tau}$, $\alpha$, $K$, $k=0$
2:An estimation of the unique solution $x_{*}$ to $Ax=f$.
3:The application definition divides the system into $k$ groups $A_{j}$ and $f_{j}$, $j\in\tau$.
4:while $k<K$ do
5:     $k=k+1$
6:     for $j\in\tau$ do, perform independent sampling $S_{j}$ for each subsystem, then update
7:         $x_{k}^{j}=x_{k-1}+\alpha\left(A_{S_{j}}^{j}\right)^{\dagger}\left(f_{S_{j}}^{j}-A_{S_{j}}^{j}x_{k-1}\right)$
8:     end for
9:     Compute $x_{k}=\sum_{j=1}^{\tau}\omega_{j}x_{k}^{j}$
10:end while
3 Orthogonal Block Kaczmarz Method Based on Preprocessing Technology
3.1 Preprocessing Technology–Sparse Reverse Cuthill-McKee Algorithm
The Kaczmarz algorithm is well known for its good convergence in linear systems, especially for sparse matrices. However, it does not exhibit fast convergence speed for all sparse matrices, particularly those with large dimensions or scattered structural distributions. Even the use of block clustering and greedy algorithms cannot effectively improve this situation. We analyze the iterative characteristics of the Kaczmarz algorithm and the matrix structure in linear systems and find that due to Kaczmarz’s orthogonal projection onto the hyperplane during each iteration, the convergence speed is closely related to the angle between the current iteration point’s hyperplane and the next projection one. It is evident that a smaller angle between these two hyperplanes leads to more iterations and slower convergence, while an angle closer to perpendicular results in fewer iterations and faster convergence.
Taking a two-dimensional linear system as an example, Figure 1 shows the iterative differences between the two cases. $H_{1}$, $H_{2}$, $H_{3}$ represent the three hyperplanes of the linear system, $x_{0}$ represents the initial point, $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ represent the results of Kaczmarz’s first, second, third, and fourth iterations, respectively. The red dots where the hyperplanes intersect represent the exact solution $x_{*}$. The left side of Figure 1 represents the iteration process with a smaller angle between $H_{1}$ and $H_{2}$, and the right side represents the iteration process when $H_{2}$ is perpendicular to $H_{3}$. It can be clearly seen that when two hyperplanes are orthogonal to each other, the Kaczmarz algorithm only needs two steps to reach the exact solution, while the non-orthogonal hyperplanes at this time are still far from the exact solution even after four steps of Kaczmarz iteration. So, we consider whether there is a method to extract rows that are orthogonal to each other in a linear system, or to process the original matrix into a matrix with rows that are orthogonal to each other? If so, the iteration time of the Kaczmarz algorithm will be greatly reduced, see [22].
@@ 3.1 Preprocessing Technology–Sparse Reverse Cuthill-McKee Algorithm
Based on the above analysis, it is considered to be cost-effective to make the original matrix into a matrix whose rows are orthogonal to each other by preprocessing the linear system. Common sparse matrix ordering algorithms include Column Count Reordering, Reverse Cuthill-McKee algorithm (RCM), Minimum Degree Reordering, Nested Dissection and Spectral Reordering, etc. [26, 27, 28, 29, 30, 31]. Each of these algorithms has its own advantages and is suitable for different types of computational tasks. Despite the various methods of ordering available, experimental comparisons have shown that the RCM algorithm usually produces superior computational results in terms of reducing bandwidth and profile of a sparse matrix. The main objective of the RCM algorithm is to minimize the maximum distance between non-zero elements and the main diagonal, thereby reducing bandwidth and greatly improving the orthogonality between matrix row blocks. Consequently, for the block Kaczmarz algorithm, the RCM algorithm is the recommended choice.
The RCM algorithm [26] was initially used for reordering symmetric matrices. It will generate a matrix with rows and columns numbered according to the node numbers. By appropriately renumbering nodes, a new matrix with much smaller matrix bandwidth can usually be generated. The specific principle is that for a symmetric sparse matrix $A$ with a given bandwidth, the RCM algorithm can provide a permutation matrix $P$ such that $\widetilde{A}=PAP^{T}$ has a smaller bandwidth, i.e., the nonzero elements of $\widetilde{A}$ are closer to the diagonal and the run-time complexity is proved to be linear in the number of nonzeros in the matrix [26].
Fortunately, the RCM algorithm also often performs equally well on asymmetric matrices (see http://www.ece.northwestern.edu/local-apps/matlabhelp/techdoc/ref/symrcm.html). Therefore, in our proposed algorithm, we first preprocess the linear system using the RCM algorithm to transform the original coefficient matrix into an approximate banded matrix. Obviously, in the approximate banded matrix, there are many rows or blocks of rows that are orthogonal to each other, see Figure 2, which improves the disadvantage of the random arrangement of the original matrix structure. Based on the above analysis, using the Kaczmarz algorithm again will greatly accelerate the convergence speed.
Figure 2:   The orthogonality representation of the blckhole matrix after being rearranged using the RCM algorithm.
@@
Note that, the run-time complexity of the RCM algorithm is proved to be linear in the number of nonzeros in the matrix [26]. Therefore, the addition of the RCM algorithm does not lead to significant time overhead. We select five matrices muu, poli3, blckhole, kim1, and wathen100 of different dimensions as examples, and count the time consumption of the RCM algorithm for preprocessing them, as shown in Table 1.
$Name$ $m$ $density\left(\%\right)$ $RCM\;preprocessing\;time\;consumption(s)$
$muu$ 7 01270127\,0127 012 $0.337$ 0.002220
$poli3$ 16 9551695516\,95516 955 $0.013$ 0.002261
$blckhole$ 2 13221322\,1322 132 $0.327$ 0.000558
$kim1$ 38 4153841538\,41538 415 $0.063$ 0.018929
$wathen100$ 30 4013040130\,40130 401 $0.051$ 0.005443
3.2 The Proposed methods